Daisha bough a car for 15000 dollars, and its value depreciated linearly. after 3 years, the value of the car was 11250 dollars. what is the amount, in dollars, of yearly depreciation?
Answer by jim_thompson5910(35256)

slope = (change in value)/(change in time) = (11250-15000)/(3-0) = -3750 = -1250

The slope is -1250 which means each year, the value of the car goes down 1250 dollars.